A rocket with a mass of 400 lbm travels at 27000 ft/sec. What is it's kinetic energy in ft-lbf
1
Expert's answer
2019-06-19T10:30:10-0400

The kinetic energy:


"K=0.5mv^2"

"m=400\ lb_m=\frac{400}{32.174}slugs"

"K=0.5\frac{400}{32.174}(27000)^2=4.5\cdot 10^9\ ft \cdot lb_f"